OpenStudy (anonymous):

-15+7(x − 1) = 5(x + 2) + 1

OpenStudy (anonymous):

-15 + 7x - 14 = 5x + 10 + 1 -29 + 7x = 5x + 11 -29 + 2x = 11 2x = 32 x = 16 I may have done this wrong though

OpenStudy (anonymous):

yes, I did it the same way you did and got it wrong.

OpenStudy (anonymous):

@Hero there's only a 1 number possible answer.... Thanks 4 your help tho! :)